The rate constant for this first‑order reaction is 0.500 s−1 at 400 degrees C.

A⟶products

How long, in seconds, would it take for the concentration of A to decrease from 0.990 M to 0.340 M?
Answer #1

We know from first order reaction

ln{[A] /[A]o} = - K * t -----(1)

K = rate constant, [A]o = Initial concentration

[A] = concentration after time t

From(1) ln(0.340 M / 0.990 M) = - K * t

ln 0.3434 = - K * t

-1.069 = - K * t

t = 1.069 / K = 1.069 / 0.500 s^-1 = 2.14 s

After 2.14 s, concentration of A to decrease from 0.990 M to 0.340 M